# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ C.1 Conceptual Framework
Following Mace (1991), I assume that the aggregate idiosyncratic shocks approach zero as $N$ gets large. Consider the situation that the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
@@
where $\sigma>0$ is the coefficient of relative risk aversion. Below, I consider a particular history of the state of the world so that the subscript $t$ can represent the dependence on the state (Hayashi et al. 1996). This simplifies the notation in equation 23 as:
@@
The aggregate consumption growth in equation(29) is assumed to reflect macroeconomic shocks in Tsukishima in the case of my empirical setting. As discussed in Section 3.4, the consumption of the THBS households indeed captures the overall macroeconomic trend. Despite this, interpreting $\alpha_{1}$ is not easy because the THBS households include the factory workers’ households in Tsukishima alone. To deal with this potential issue, I use a two-way fixed-effects model instead of the first-difference model in the spirit of Cochrane (1991). Another potential issue is that $\epsilon_{i,t}$ may include the household-specific time-varying preference shock, which might be correlated with the idiosyncratic income shock. Family size influences the household’s preference because the marginal utility of consumption increases with family size (Jappelli and Pistaferri 2017). As explained, the family size had been fairly stable during the sample period in the THBS households, meaning that the preference shifts had rarely occurred (Section 3.4). To be conservative, however, I include the number of family members as a control variable for representing potential preference shock.
@@
In order to test the risk-coping mechanisms, I consider the following representation of individual consumption in the spirit of Fufchamps and Lund (2003).
@@
where $y_{i}^{P}$ is permanent income and $y_{i,t}^{T}$ is transitory income. $r_{i,t}=\bm{z}^{\prime}_{i,t}\bm{1}$ is a linear combination of the net temporary incomes expressed using a $k\times 1$ net income vector ($\bm{z}$) and a vector of ones. Substituting equation 31 into equation 26 and taking the first-difference removes the time-constant components, including the permanent income. It follows that my empirical specification can be characterized as follows:
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.112112112In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue rises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.2 Raw Relationship between Income and Expenditures
Generally, rent was paid once per month. This means that the semi-monthly dataset fails to capture the variations in rent payment because it includes some semi-monthly cells with zero values. In addition, the water bills are included in the rent subcategory but not in the utility subcategory. This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell. Figure C.1c indeed shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments. These cells taking large LD values substantially increased the slope, as shown in Figure C.2c.
@@
Sources: Created by the author using the THBS sample.
@@
Notes: Figure C.2a illustrates the relationship between the log-differences in semi-monthly income and expenditure. Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
Sources: Created by the author using the THBS sample.
@@
Sources: Created by the author using the THBS sample.
@@
Notes: Figure C.4a illustrates the relationship between the log-differences in semi-monthly income and expenditure. Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
@@ C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, greengrocers, and fuels (firewood and charcoal) by scale of business. In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%)”. For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash”, 78.4% for “Credit purchase”, and 1.3% for “Monthly installment”.
@@
This table summarizes the statistics on the sales in the retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
@@ C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
Table C.2 summarizes the use of credit purchase and the losses on the credit sales documented in the 1930 Survey Report. Following the report’s style, the item categories are divided into 6 categories: white rice, other grains, vegetables, fruit, fish, and fuel. The total number of retailers in each category is 157, 7, 67, 16, 46, and 64, respectively. As explained in Section 5, most of them used credit purchases as their selling method. Importantly, among the retailers using credit purchases, almost all of them suffered losses in sales. For example, 99% of white rice retailers experienced any losses. The minimum value of this share is still 97% for the fuel retailers. This means that credit purchase was inherently a loss-making sales method for the retailers.
The statistics in Table C.2 also indicate that many cases have unknown loss amounts. Overall, however, most of the cases seem to be grouped into bins of less than 5% of losses. For example, 87% (97 retailers) of the 111 available rice retailers suffered losses amounting to 5% of sales. Similarly, 65% (32 retailers) of the 49 available fuel retailers are grouped into the same loss category. Systematic statistics are unavailable to indicate the percentage of total profits corresponding to these losses. However, it is worth mentioning that a large part of losses was capped by 5% of total sales in all these different retailers. In other words, this implies that while the retailers could not avoid losses, they might have set prices that anticipated predicted losses.
@@
Notes: This table summarizes the statistics on the sales methods and losses on sales under credit purchases among the retailers measured in Tokyo City surveyed in 1930. Source: Tokyo City Office 1932, pp. 454–469.
@@
  • (6) Choi, In “Unit root tests for panel data”. Journal of International Money and Finance 20, no. 2 (2001): 249–272.
@@
  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behaviour before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
@@
  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference”. Journal of Political Economy 119, no.5, (2011): 925–958.
@@
  • (66) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, volume. 2, Pubic pawnshops) [in Japanese]. Tokyo Institute for Municipal Research, 1926.
@@
  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-2).
  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-3).
  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-4).
@@
  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-12).
  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-13).
  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-14).
  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-15).

### v10

@@ C.1 Conceptual Framework
Following Mace (1991), I assume that the aggregate idiosyncratic shocks approach zero as $N$ gets large. Consider the situation in which the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
@@
where $\sigma>0$ is the coefficient of relative risk aversion. Below, I consider a particular history of the state of the world so that the subscript $t$ can represent dependence on the state (Hayashi et al. 1996). This simplifies the notation in equation 23 as:
@@
The aggregate consumption growth in equation(29) is assumed to reflect macroeconomic shocks in Tsukishima in the case of my empirical setting. As discussed in Section 3.4, the consumption of the THBS households captures the overall macroeconomic trend. Despite this, interpreting $\alpha_{1}$ is difficult because the THBS households include the factory workers’ households in Tsukishima alone. To manage this potential issue, I use a two-way fixed-effects model instead of the first-difference model, following Cochrane (1991). Another potential issue is that $\epsilon_{i,t}$ may include household-specific time-varying preference shock, which might be correlated with idiosyncratic income shock. Family size influences the household’s preference because the marginal utility of consumption increases with family size (Jappelli and Pistaferri 2017). As explained, family size was stable during the sample period in the THBS households, meaning that the preference shifts rarely occurred (Section 3.4). To be conservative, however, I include the number of family members as a control variable for representing potential preference shock.
@@
To test the risk-coping mechanisms, I consider the following representation of individual consumption following Fufchamps and Lund (2003).
@@
where $y_{i}^{P}$ is permanent income and $y_{i,t}^{T}$ is transitory income. $r_{i,t}=\bm{z}^{\prime}_{i,t}\bm{1}$ is a linear combination of the net temporary incomes expressed using a $k\times 1$ net income vector ($\bm{z}$) and a vector of ones. Substituting equation 31 into equation 26 and taking the first-difference removes the time-constant components, including permanent income. It follows that my empirical specification can be characterized as follows:
@@
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.2 Raw Relationship between Income and Expenditures
Generally, rent was paid once per month. This means that the semi-monthly dataset fails to capture the variations in rent payment because it includes some semi-monthly cells with zero values. In addition, the water bills are included in the rent subcategory but not in the utility subcategory. This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell. Figure C.1c shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments. These cells taking large LD values substantially increase the slope, as shown in Figure C.2c.
@@
Source: Created by the author using the THBS sample.
@@
Notes: Figure C.2a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure. Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
Source: Created by the author using the THBS sample.
@@
Source: Created by the author using the THBS sample.
@@
Notes: Figure C.4a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure. Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
@@ C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, fruit and vegetable, and fuels (firewood and charcoal) by scale of business. In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%).” For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash,” 78.4% for “Credit purchase,” and 1.3% for “Monthly installment.”
@@
This table summarizes the statistics on the sales in the retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
@@ C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
Table C.2 summarizes the use of credit purchase and the losses on the credit sales documented in the 1930 Survey Report. Following the report’s style, the item categories are divided into six categories: white rice, other grains, vegetables, fruit, fish, and fuel. The total number of retailers in each category is $157$, $7$, $67$, $16$, $46$, and $64$, respectively. As explained in Section 5, most of them used credit purchases as their selling method. Importantly, among the retailers using credit purchases, almost all of them suffered losses in sales. For example, $99$% of white rice retailers experienced losses. The minimum value of this share is still $97$% for the fuel retailers. This means that credit purchase was a loss-making sales method for the retailers.
The statistics in Table C.2 also indicate that many cases have unknown loss amounts. Overall, however, most of the cases seem to be grouped into bins of less than $5$% of losses. For example, $87$% ($97$ retailers) of the $111$ available rice retailers suffered losses amounting to $5$% of sales. Similarly, $65$% ($32$ retailers) of the $49$ available fuel retailers are grouped into the same loss category. Systematic statistics are unavailable to indicate the percentage of total profits corresponding to these losses. However, it is worth mentioning that a large part of losses was capped by $5$% of total sales in all these different retailers. This implies that while the retailers could not avoid losses, they might have set prices that anticipated losses.
@@
Note: This table summarizes the statistics on the sales methods and losses on sales under credit purchases among the retailers measured in Tokyo City surveyed in 1930. Source: Tokyo City Office 1932, pp. 454–469.
@@
  • (6) Choi, In “Unit root tests for panel data.” Journal of International Money and Finance 20, no. 2 (2001): 249–272.
@@
  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behavior before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
@@
  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference.” Journal of Political Economy 119, no.5, (2011): 925–958.
@@
  • (66) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, Volume. 2, Public pawnshops) [in Japanese]. Tokyo Institute for Municipal Research, 1926.
@@
  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7–2).
  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7–3).
  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7–4).
@@
  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–12).
  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–13).
  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–14).
  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–15).